f 25.5 L of oxygen are heated from 50oC to 150oC at constant pressure, what is the new volume of oxygen?

0.0514 L

19.5 L

33.4 L

0.03 L
Answer #1

Boyles law - This law states that at constant temperature, The pressure of a fixed amount of gas is inversely proportional vTvat] ب- آه| د- اه | V= 2S2 T = S6 + ا 323 = 3 Tz =1507273=423k ریه کر کے - - - یہ کا 23:38 23 33.3717 ۷- 33 با 330 = rate pl